From nch@roe.ac.uk Mon Jul 28 15:15:16 2008 Date: Mon, 28 Jul 2008 10:02:10 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 25/07/08 Minutes of WFAU VDFS Science Archive meeting: 25th July 2008 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, RPB, ETWS, RSC, NJC, MAR, RGM, LGR Apologies: JPE, JMS, AL, BCW, MSH, PMW DoNM: N.B. 11am, Friday 1st August 2008 in the Vista Hut Actions discharged this week: ----------------------------- ACTION: RPB to finish off rearranging the SSA. Discharged; also thanks to MAR and ETWS for sorting out the webdocs and deployment of the new all-sky service (which stayed online for all of one day... see Hardware below). RPB noted that the SSA on amenhotep is now backed up to tape, though the database has not been scrubbbed yet. ACTION: MSH to check out thoth and reboot scheduled Sunday evening. Discharged Actions partly discharged but continuing: ----------------------------------------- None this week Actions carried forward from 18/07/08 meeting: ---------------------------------------------- ACTION: RPB to formulate backup plan arising from meeting Continues; currently wholesale new backups running; formal policy to be documented. NCH noted that he has made a TWiki page with information on the legacy SuperCOS dataset backups, and this could be used as a model. Specific points and new actions: -------------------------------- Project management: Nothing new this week. WFCAM & VISTA updates: PSB has noted that CASU have some new VISTA frames and may be able to provide WFAU with some processed data products soon. Comments and issues arising from CASU fortnightly minutes: No new minutes this week Networking: ETWS noted that UKLight has been a bit up and down recently and this has affected routine transfer ops; coupled with problems with the NAS box (see below) general ingest ops are a few days behind schedule, but nonetheless RPB noted that just two weeks of 08A data remain to be transfered. Earlier in the week, PSB informed us of the latest addition to the CASU VLAN - a Sun X4500 which will be the first mass storage for VISTA. This has been successfully pinged from djedefre at the WFAU end. WSA Operations: RPB noted a lot of problems with takelot this week: "On Friday it became completely unresponsive and Mark had to switch it off and on again. On Tuesday night it stopped talking to the network, although the console was perfectly responsive. I rebooted and all was well, although all network shares had to be remounted by IT Support. Same thing happened on Wednesday night. Thursday morning it again refused to talk to the network. This time, however, I changed the configuration so that it used its second onboard ethernet card. Network shares came back up with no input from IT Support, and takelot has been up ever since (though since most servers have been switched off due to the warm weather, that may be another factor). Copying data up from Cambridge continues apace, though majorly hampered by takelot outages. All of May's data has now copied, and June's started copying this morning. Some of May's data has been ingested, though this again was hampered by takelot's problems. Lots of starting, stopping, and checking meant that this took up much more time than it had any right to." Hardware Main issue this week has been the nice warm weather and consequent overheating (again) in server room C1. ITSG requested WFAU shutdown non-essential servers yesterday afternoon. All public WSA and one operations server remain up so users should not be affected by the problem, but unfortunately the DB server hosting the new all-sky SSA had to be shutdown one day after deploying the service. NCH suggested that we wait until Monday next week before starting to switch everything back on again as the weather forecast is good for the next few days. NAS box takelot has been misbehaving this week; RPB suspects a dodgy network interface card. Eclipse have been called in and will examine the unit early next week. MSH has distributed a comprehensive SAN plan as regards new, high capacity database storage: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SANStorageAnalysis (for internal consumption only). NCH noted that apart from the increase in IO bandwidth, a critical issue is the ability to attach read-only DBs to multiple instances of SQL server, and hence we need to test out attaching an existing database in such a mode using our existing instances of SQL 2005. ACTION: RPB to test out attaching a normal database as an MS SQL "reporting database" in readonly mode on thutmose/hatshepsut Software: MAR noted "Switched over SSA interface/webpages to all-sky DB on hatshepsut. Also started looking at implementing SSA interface on horus. At user request modded WSA UI code to allow getImage linking of SQL results for GPS pre DR3 schema." NJC reported: "This week NJC made improvements to the production of synoptic tables, including more control and testing structures, adding a motion model parameter into the varFrameSetInfo and tested this out on some more non-surveys. NJC also added more figures and improved the text on the synoptic paper that he is preparing. After testing whether it is possible to curate a programme that has some single epoch filters configured to be used as a "shallow" survey setup and some multi-epoch filters configured in a "deep" survey setup, NJC concludes that it is safer to configure all filters to be "deep" if one or more is multi-epoch." ETWS and RSC both noted that there are software compatibility issues with 64-bit Linux. CU3 currently goes completely bananas, and CU7 is not much better. The suspicion is that the problem lies with our own software, and some time will need to be devoted to checking this out. NJC noted that we should check running of the toolkit codes provided by CASU also. Survey Data Release: SJW has reminded UKIDSS survey heads to keep up with the eyeball QC in 08A (action NCH...). Non-survey Data Release: Nothing new to report this week. Astrogrid deployment & Data Analysis services LGR noted that the optical/IR pixel analysis tool is now ready for testing. Miscellaneous: RPB suggests moving next week's meeting back an hour to let people watch the partial eclipse on Friday. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================